OpenStudy (anonymous):

Given cos x = -4/7, find cos 2x (assume x is in Quadrant III)

terenzreignz (terenzreignz):

Okay... if x is in quadrant 3, where is 2x?

OpenStudy (anonymous):

no clue. Thats what i'd like to know as well..

terenzreignz (terenzreignz):

Do you work with degrees or radians?

OpenStudy (anonymous):

radians

terenzreignz (terenzreignz):

Okay, if x is in the third quadrant (Q3) then \[\huge \pi < x < \frac{3\pi}{2}\] That said, where is 2x?